OpenStudy (anonymous):

Melody spent 20% of her pocket money on junk food and 40% of the remainder on the make up. she deposited the remaining $24 in her bank account. How much did she spend on make up?

OpenStudy (anonymous):

percentage of money left with her = 100 - (20 +(40*(100-20)) =100- 52= 48 thus she has spent (32/100) (24/48) *100 = 16

OpenStudy (anonymous):

yes, $ 16

OpenStudy (anonymous):

i still dont understand.